OpenStudy (anonymous):

derivative of ln8x^2

OpenStudy (freckles):

\[\frac{d}{dx} \ln(8x^2)=\frac{(8x^2)'}{8x^2}\]

OpenStudy (anonymous):

So it would be ln8x^2 multiply 1/8x^2

OpenStudy (freckles):

No

OpenStudy (freckles):

let me rewrite what I said a different way \[\frac{d}{dx}(\ln(8x^2))=\frac{\frac{d}{dx}(8x^2)}{8x^2}\]

OpenStudy (anonymous):

16x/8x^2

OpenStudy (anonymous):

which is 8/x

OpenStudy (anonymous):

Arent you leaving out the ln

OpenStudy (freckles):

2/x.. If you meant a different function from the one I wrote maybe